### v2

@@ Abstract
We introduce MHNpath, a machine learning-driven retrosynthetic tool designed for computer-aided synthesis planning. Leveraging modern Hopfield networks and novel comparative metrics, MHNpath efficiently prioritizes reaction templates, improving the scalability and accuracy of retrosynthetic predictions. The tool incorporates a tunable scoring system that allows users to prioritize pathways based on cost, reaction temperature, and toxicity, thereby facilitating the design of greener and cost-effective reaction routes. We demonstrate its effectiveness through case studies involving complex molecules from ChemByDesign, showcasing its ability to predict novel synthetic and enzymatic pathways. Furthermore, we benchmark MHNpath against existing frameworks, replicating experimentally validated “gold-standard” pathways from PaRoutes. Our case studies reveal that the tool can generate shorter, cheaper, moderate-temperature routes employing green solvents, as exemplified by compounds such as dronabinol, arformoterol, and lupinine.
Keywords Retrosynthetic Planning  $\cdot$ Modern Hopfield Networks  $\cdot$ Computer-aided Synthesis Planning
@@ 1 Introduction
The integration of machine learning (ML) into chemical synthesis has transformed the field of computer-aided synthesis planning (CASP), providing chemists with powerful tools to design and execute synthetic routes more efficiently. Traditional retrosynthetic analysis, which relies heavily on expert intuition and experience 1, 2, is increasingly being augmented by data-driven approaches 3 to break down complex molecules into simpler precursors using ML techniques to predict viable synthetic pathways 4, 5. These advancements are particularly crucial in addressing the growing complexity of modern synthetic challenges, where exploring vast chemical spaces catalyzes the discovery of novel molecules. Examples of such challenges include the synthesis of complex natural products with multiple stereocenters 6 and the need for highly selective functional group transformations in pharmaceutical development 7.
@@
In CASP, clean chemical data is the foundation for designing synthetic routes. This data fuels a predictive model that finds precursors while an efficient search algorithm navigates the expansive chemical space to propose feasible retrosynthetic pathways. Complementing these elements is an integrated scoring system that evaluates routes based on multiple criteria, such as cost and reaction temperature. Together, these components establish a cohesive framework that underpins modern computer-aided synthesis planning.
State-of-the-art machine learning models for CASP can be categorized into template-based and template-free approaches; templates are generalized representations that encapsulate the core chemical transformation patterns inherent to each reaction.
@@
Recent advancements in ML, including the development of deep neural networks and transformer-based models, have significantly enhanced the predictive capabilities of template-free methods in CASP 8, 9, 10, 11, 12. Transformer models have emerged as a powerful tool in retrosynthetic prediction due to their ability to handle sequence-to-sequence tasks effectively. The Molecular Transformer model 9, and its variants 10, 11, 12, for instance, has demonstrated remarkable accuracy in predicting both reactants and reaction conditions by treating retrosynthesis as a language translation problem from products to reactants. These models leverage self-attention mechanisms to capture complex dependencies within chemical reactions, making them highly effective for single-step retrosynthesis 13. Some graph-based methods 14, 15 predict the product from reactants in an auto-regressive fashion, while others 16, 17, 18 predict the product by anticipating the final states of bonds or electrons. However, their reliance on extensive labeled reaction data, computational resources, and single-step focus presents a significant limitation.
@@
Rule-based approaches have been fundamental in CASP 19. These approaches apply predefined reaction rules to fragment target molecules into simpler structures. Rule-based methods are interpretable and align well with chemical intuition. Segler et al.20 and Coley et al.21 were among the first ones to use automatically extracted rules to predict outcomes of organic reactions using neural networks. Chen et al. 22 proposed a chemistry-motivated graph neural network that uses templates to describe the net changes in electron configuration between the reactants and products. Recent advancements, such as the use of Modern Hopfield Networks for template prediction, have improved the performance of rule-based CASP tools 23. Hopfield networks allow for the efficient prioritization of reaction templates by leveraging associative memory mechanisms, significantly improving the speed and reliability of retrosynthetic planning. Despite these improvements, rule-based methods often struggle with the rigidity of predefined rules, which can limit their ability to propose novel synthetic routes or adapt to emerging reaction types.
@@
Retrosynthesis has traditionally been treated as a tree search problem, where each step involves searching for chemically feasible precursors to derive the product molecule. Initial works used greedy search 5, but more recent efforts have adopted Monte Carlo tree search 24 and A-like algorithms 25, 26. These search methodologies attempt to optimize the exploration of synthetic routes by balancing the exploitation of high-confidence reaction pathways with the exploration of novel or underutilized transformations. However, their effectiveness is contingent on the quality and comprehensiveness of the reaction databases they utilize.
@@
In addition to these methodologies, the integration of biocatalysis in retrosynthetic planning has gained attention for its potential to enhance sustainability. Biocatalysis leverages enzymatic transformations to achieve highly selective reactions under mild conditions, reducing the environmental footprint of chemical synthesis 27. Incorporating enzymatic steps into synthetic pathways is particularly advantageous for pharmaceutical and fine chemical industries, where stringent purity and selectivity requirements are critical 28, 29. Tools like RetroBioCat 27 have demonstrated the ability to design biocatalytic cascades that complement synthetic routes by introducing enzymatic steps that are often more selective and environmentally friendly. However, even these specialized tools are limited by the availability of comprehensive enzymatic reaction data and their focus on narrow reaction types.
In this work, we introduce MHNpath, an ML-driven retrosynthetic tool designed to provide chemists with a comprehensive platform for exploring greener, cost-effective, and moderate-temperature reaction pathways. The primary contributions of this study are threefold: (1) the development of a robust predictive model that efficiently prioritizes reaction templates using Modern Hopfield Networks 30 and the introduction of two new metrics for a more comprehensive comparison of template prioritizers; (2) the implementation of a user-tunable scoring system that allows users to prioritize pathways based on criteria such as precursor chemicals’ cost, reaction temperature, and toxicity of solvents; and (3) the demonstration of MHNpath’s effectiveness through case studies involving the synthesis of complex molecules from PaRoutes 31 and ChemByDesign 32.
The significant components of MHNpath include processed enzymatic and synthetic datasets with algorithmically extracted templates, a modified Modern Hopfield Network-based architecture for template prioritization, a user-tunable scoring strategy, and a global greedy tree search algorithm. By addressing the limitations of existing CASP tools - such as their heavy data requirements, single-step focus, and rigidity in rule-based methods - MHNpath aims to facilitate the discovery of innovative synthetic routes and promote sustainable practices in chemical synthesis.
The following sections will detail the methodology behind MHNpath, including the data processing pipeline, the implementation of the Modern Hopfield Network-based template prioritizer, and the tree search methodology employed for retrosynthetic exploration, as outlined in Figure 1. Furthermore, we will present results comparing MHNpath against existing state-of-the-art CASP tools and provide case studies demonstrating its practical applications.
@@ Methods
The MHNpath framework is presented in Figure 1. A modified Modern Hopfield Networks 30 based architecture by Seidl et al. 23 is used. Each prioritizer accepts a target molecule as input and predicts a ranked list of the most applicable reaction templates, which are applied to the target molecule to get precursors (nodes). The individual scores for precursor cost, reaction temperature, and solvent toxicity are calculated; these are summed up using user-tunable weights. The precursors are further explored in a recursive manner using a global greedy tree search algorithm, where the best-scoring unexplored node from the entire tree is explored first. Figure 1(b) gives an example of the multi-modal-guided global greedy tree search approach.
Data
@@
Our study leveraged two comprehensive datasets to train and evaluate the MHNpath framework. The first dataset, focused on enzymatic reactions, was compiled from the BKMS dataset 33 and the Rhea database 34, initially containing about 68,000 reactions. Following a rigorous preprocessing procedure, which involved removing duplicates, invalid SMILES, and unbalanced reactions, we obtained 35,289 unique and valid reaction SMILES. We extracted 17,047 reaction templates, which are generalized representations that encapsulate the core chemical transformation patterns inherent to each reaction, from this refined set using the RDKit library 35.
The second dataset, focused on synthetic organic reactions, was sourced from the USPTO dataset 36. Starting with 1,808,937 reactions, we applied a series of preprocessing steps, which resulted in 1,693,109 refined reactions. From this set, we extracted 301,242 reaction templates. We employed a scaffold split technique to optimize computational efficiency and model performance using Datamol 37 to divide the synthetic reactions into five equally sized subsets. This division ensured a balanced distribution of chemical diversity, a crucial factor for training a robust model capable of capturing a wide variety of synthetic reactions.
The rationale behind splitting the dataset into five subsets stems from the need to manage computational resources effectively while maintaining a high diversity level in the training process. Large datasets, such as those containing millions of reactions, can be computationally intensive and may cause significant memory overheads when processed as a whole. By splitting the data into smaller subsets, we reduce the per-subset size and ensure that each subset can be handled efficiently in terms of memory and processing time. This partitioning strategy ensures that we can still explore the broad chemical space of synthetic reactions without overwhelming the computational resources available.
Furthermore, splitting the data allows us to optimize training in a more controlled manner. Each subset is drawn from the entire chemical space, which ensures that every model trained on a subset is exposed to a representative sample of the overall data. A single synthetic template prioritizer achieves a top 1 accuracy of 36.6% on its own test set compared to a top 1 accuracy of 42.2% obtained by the ensemble method on the whole test set. This results in a more robust overall model when the outputs of all five models are combined using an ensemble approach.
To further enhance the predictive power of our framework, we trained five smaller synthetic template prioritizers using these subsets and employed an ensemble approach to make predictions. The ensemble method aggregates the predictions from each of the five models, which helps to improve the accuracy and robustness of the framework. By leveraging the predictions from various models focusing on different data portions, the ensemble approach reduces the risk of overfitting to a specific subset and increases the model’s generalizability.
The choice of these datasets is important as they provide a comprehensive representation of both enzymatic and synthetic reactions, which are essential for developing robust predictive models in our study.
Model
@@
We employ the Modern Hopfield Networks (MHN) 30 based template prioritization architecture for our study. This architecture, initially introduced by Seidl et al.23, consists of three main components: a molecule encoder, a reaction template encoder, and one or more stacked or parallel Hopfield layers.
@@
The molecule encoder function, denoted as $h_{wm}(m)$, learns a relevant representation for the input molecule $m$. We utilize a fingerprint-based approach, particularly the extended connectivity fingerprint (ECFP), coupled with a fully connected neural network with weights $w$. This encoder maps a molecule to a representation $m_{h}=h_{w}^{m}(m)$ of dimension $d_{m}$. The molecule encoder is crucial for transforming molecular structures into a format suitable for the model.
@@
Similarly, the reaction template encoder function, denoted as $h_{v}^{t}(T)$, learns relevant representations of reaction templates $T$. We also employ a fully connected neural network with RDKit 35 template fingerprints as input. The resulting vectors are concatenated column-wise into a matrix $T_{h}=h_{v}^{t}(T)$ with shape $(d_{t},K)$. This encoder is essential for transforming reaction templates into a comparable format.
Both encoders are equipped with Xavier Initialization 38 and Batch Normalization 39. Xavier Initialization significantly improves the model’s performance by initializing weights to prevent vanishing/exploding gradients.
The model’s core consists of one or more stacked or parallel Hopfield layers, denoted as $g(.,.)$, which associate a molecule with all templates in memory. Hopfield layers incorporate layer normalization for $\xi$ and $X$, a hyperparameter, and a scaling parameter $\beta$. The update rule of the Hopfield layer is described by the following equation, where the updated representation of the product molecule $\xi_{new}$ and the vector of associations $p$ are obtained:
@@
$\xi_{new}$ $=Xp=X\cdot\text{softmax}(\beta X^{T}\xi)$ (1)
For the loss function and optimization, given a training pair $(m,t)$ and the set of all templates $T$, the model aims to assign a high probability to $t$ based on $m$ and $T$. The negative log-likelihood is employed as the loss function, with parameters adjusted using stochastic gradient descent via the AdamW optimizer.
To prevent overfitting, regularization techniques such as dropout and L2 regularization are applied, along with a post-processing step involving a fingerprint-based substructure screen to filter out non-applicable templates.
We train two sets of models on our datasets: the Synthetic Template Prioritizers, consisting of five models, and the Enzymatic Template Prioritizer. These models take a target molecule as input and rank the most applicable rules within the collected dataset. Given the same inputs, the outputs from the five Synthetic Template Prioritizers are taken individually, and a final ranking of the templates is collated using the highest predicted score.
Furthermore, to fine-tune the models for optimal performance, we employed a manual Bayesian search method for hyperparameter tuning 40. This approach allowed us to systematically explore the hyperparameter space, identifying configurations that yielded superior results across various evaluation metrics.
Searching

### v3

@@ Abstract
We introduce MHNpath, a machine learning-driven retrosynthetic tool designed for computer-aided synthesis planning. Leveraging modern Hopfield networks and novel comparative metrics, MHNpath efficiently prioritizes reaction templates, improving the scalability and accuracy of retrosynthetic predictions. The tool incorporates a tunable scoring system that allows users to prioritize pathways based on cost, reaction temperature, and toxicity, thereby facilitating the design of greener and cost-effective reaction routes. We demonstrate its effectiveness through case studies involving complex molecules from ChemByDesign, showcasing its ability to predict novel synthetic and enzymatic pathways. Furthermore, we benchmark MHNpath against existing frameworks using the PaRoutes dataset, achieving a solution rate of 85.4% and replicating 69.2% of experimentally validated “gold-standard” pathways. Our case studies show that the tool can generate shorter, cheaper, moderate-temperature routes employing green solvents, as exemplified by molecules such as dronabinol, arformoterol, and lupinine.
@@ 1    Introduction
The integration of machine learning (ML) into chemical synthesis has transformed the field of computer-aided synthesis planning (CASP), providing chemists with powerful tools to design and execute synthetic routes more efficiently. Traditional retrosynthetic analysis, which relies heavily on expert intuition and experience Szymkuc2016, PENSAK1977, is increasingly being augmented by data-driven approaches D4DD00120F to break down complex molecules into simpler precursors using ML techniques to predict viable synthetic pathways szymkuc2016computer, levin2022merging. These advancements are particularly crucial in addressing the growing complexity of modern synthetic challenges, where exploring vast chemical spaces catalyzes the discovery of novel molecules. Examples of such challenges include the synthesis of complex natural products with multiple stereocenters nicolaou2008molecules and the need for highly selective functional group transformations in pharmaceutical development brown2016analysis.
@@
In CASP, clean chemical data is the foundation for computationally designing synthetic routes. This data enables training of a predictive model that finds precursors while an efficient search algorithm navigates the expansive chemical space to propose feasible retrosynthetic pathways. State-of-the-art ML models for CASP can be categorized into template-based and template-free approaches where templates are generalized representations that encapsulate the core chemical transformation patterns inherent to each reaction. Complementing the ML model and search algorithm is an integrated scoring system that evaluates routes based on multiple criteria, such as cost and reaction temperature. Together, these three components establish a cohesive framework that underpins modern computer-aided synthesis planning.
@@
Recent advancements in ML, including the development of deep neural networks and transformer-based models, have significantly enhanced the predictive capabilities of template-free methods in CASP hart1968formal, schwaller2019molecular, D1SC02362D, tetko2020state, Irwin_2022. Transformer models have emerged as a powerful tool in retrosynthetic prediction due to their ability to handle sequence-to-sequence tasks effectively D5DD00153F. The Molecular Transformer model schwaller2019molecular, and its variants D1SC02362D, tetko2020state, Irwin_2022, for instance, has demonstrated remarkable accuracy in predicting both reactants and reaction conditions by treating retrosynthesis as a language translation problem from products to reactants. These models leverage self-attention mechanisms to capture complex dependencies within chemical reactions, making them highly effective for single-step retrosynthesis WANG2021129845, andronov2025specbeam. Some graph-based methods bradshaw2018a, sacha2021molecule predict the products from reactants in an auto-regressive fashion, while others C8SC04228D, qian2020integrating, pmlr-v139-bi21a predict the products by anticipating the final states of bonds or electrons. However, the reliance of template-free approaches on extensive labeled reaction data, computational resources, and single-step focus has been a significant limitation in CASP.
@@
On the other hand, template-based approaches have also been fundamental in CASP.corey1985computer These approaches directly apply predefined reaction rules to fragment target molecules into simpler structures, making them more interpretable and chemically intuitive compared to template-free approaches. Segler et al.segler2017neural and Coley et al.coley2017prediction were among the first ones to use automatically extracted rules to predict outcomes of organic reactions using neural networks while Zhang et al. zhang2025groupretro extend it by also continually evolving them and Rho et al. roh2025higherlevel give a global outlook by effectively abstracting the detailed substructures. Chen et al. chen2022generalized proposed a chemistry-motivated graph neural network that uses templates to describe the net changes in electron configuration between the reactants and products. Recent advancements, such as the use of Modern Hopfield Networks for template prediction, have improved the performance of template-based CASP tools seidl2022improving. Modern Hopfield Networks allow for the efficient prioritization of reaction templates by leveraging associative memory mechanisms, significantly improving the speed and reliability of retrosynthetic planning. Nevertheless, template-based methods often struggle with the rigidity of predefined rules, analyzed by Choe et al. choe2025crosstalk, which can limit their ability to propose novel synthetic routes or adapt to emerging reaction types.
@@
Computational retrosynthesis has traditionally been treated as a tree search problem, where each step involves searching for chemically feasible precursors to derive the product molecule. Initial works have used greedy search levin2022merging, but more recent efforts have adopted Monte Carlo tree search D0SC04184J and A-like algorithms chen2020retro, yu2024doubleendedsynthesisplanninggoalconstrained. These search methodologies attempt to optimize the exploration of synthetic routes by balancing the exploitation of high-confidence reaction pathways with the exploration of novel or underutilized transformations. However, their effectiveness is contingent on the quality and comprehensiveness of the reaction databases they utilize.
@@
In addition to the different components of CASP, the integration of biocatalysis in computational retrosynthetic planning has gained attention for its potential to enhance sustainability. Biocatalysis leverages enzymatic transformations to achieve highly selective reactions under mild conditions, reducing the environmental footprint of chemical synthesis finnigan2021retrobiocat. Incorporating enzymatic steps into synthetic pathways is particularly advantageous for pharmaceutical and fine chemical industries, where stringent purity and selectivity requirements are critical B713736M, Horvath2007. Tools like RetroBioCat finnigan2021retrobiocat have demonstrated the ability to design biocatalytic cascades that complement synthetic routes by introducing enzymatic steps that are often more selective and environmentally friendly. However, even these specialized tools are limited by the availability of comprehensive enzymatic reaction data and their focus on narrow reaction types.
In this work, we introduce MHNpath, an ML-driven retrosynthetic tool designed to help chemists explore greener, cost-effective, and moderate-temperature synthetic pathways. First, we develop a robust template-prioritization model based on Modern Hopfield Networks ramsauer2020hopfield, enhanced with Xavier initialization, dropout, and L2 regularization to improve stability and generalization. We additionally introduce two new evaluation metrics that enable a more comprehensive comparison of template prioritizers. Second, we implement a user-tunable scoring system that integrates practical constraints and include precursor cost, reaction temperature, and solvent toxicity, allowing chemists to steer the search toward more sustainable and operationally feasible routes. Third, we demonstrate MHNpath’s effectiveness through case studies from PaRoutes D2DD00015F and ChemByDesign draghici2012chemistry, showcasing its ability to identify viable, interpretable multi-step pathways. Together with its curated enzymatic and synthetic template datasets and a global greedy tree-search algorithm, MHNpath provides a flexible and data-efficient framework that addresses the limitations of existing CASP tools, particularly their reliance on extensive training and rigid template-based systems. The following sections will detail the methodology behind MHNpath, including the data processing pipeline (Section 2.1), the implementation of the Modern Hopfield Network-based template prioritizer (Section 2.2), and the tree search methodology and custom scoring employed for retrosynthetic exploration (Sections 2.3 & 2.4), as outlined in Figure˜1. Furthermore, we will present results (Section 3) comparing MHNpath against existing CASP tools and provide case studies demonstrating its practical applications.
@@ 2    Methods
The MHNpath framework is presented in Figure˜1. A modified Modern Hopfield Network ramsauer2020hopfield based architecture by Seidl et al. seidl2022improving is adapted. Each prioritizer (also referred to as model) accepts a target molecule as input and predicts a ranked list of the most applicable reaction templates, which are applied to the target molecule to get precursors (nodes). The individual scores for precursor cost, reaction temperature, and solvent toxicity are calculated and summed up using user-tunable weights. The precursors are further explored in a recursive manner using a global greedy tree search algorithm where the best-scoring unexplored node from the entire tree is explored first. Figure˜1(b) gives an example of the multi-modal-guided global greedy tree search approach.
2.1   Dataset processing
@@
Our study leveraged two comprehensive training datasets to develop the predictive component of the MHNpath framework. Additionally, specific external datasets were curated for benchmarking and evaluation tasks. In this context, we selected training data that provided a comprehensive representation of both enzymatic and synthetic reactions. The training dataset of enzymatic reactions was compiled from the BKMSlang2011bkm dataset and the Rheabansal2022rhea database, initially containing about 68,000 reactions. Following a rigorous preprocessing procedure that involved removing duplicates, invalid SMILES representations, and unbalanced reactions, we obtained 35,289 unique and valid reaction SMILES. From this refined enzymatic subset, we extracted 17,047 reaction templates using the RDKitLandrum2016RDKit2016_09_4 library. This subset was used to train a single enzymatic template prioritizer, utilizing an 80:10:10 random split for training, validation, and testing, respectively.
We also used a dataset of synthetic organic reactions for training which was sourced from the USPTOLowe2017 dataset. Starting with 1,808,937 reactions and applying a series of preprocessing steps resulted in 1,693,109 refined reactions, from which 301,242 reaction templates were extracted in the synthetic subset. To manage computational resources and memory overheads due to the large number of data points in the synthetic subset, we employed a scaffold split technique using Datamol hadrien_mary_2024_10535844 to ensure high chemical diversity while dividing the subset further into five equally sized subsets. This partitioning strategy allowed us to train five distinct synthetic template prioritizers (one per divided synthetic subset). Each divided subset underwent an 80:10:10 random split for model development. During inference, the predictions from these five models are aggregated using an ensemble approach. This strategy significantly enhances robustness with a single synthetic template prioritizer achieving a top-1 accuracy of 36.6% on its own test set and the ensemble method reaching a top-1 accuracy of 42.2% on the combined test set. In this context, top-1 accuracy denotes the model’s ability to accurately prioritize the correct ground-truth template as the number one rank in the prediction list.
Once the predictive components of MHNpath were developed, we rigorously assessed the framework’s performance in Sections 3.2 and 3.3 using three tasks as described below. These tasks required curating distinct datasets from the literature and provide unbiased performance metrics for systems not present in the training set.
  • PaRoutes benchmark: To evaluate pathway reconstruction against patent literature, we utilized the PaRoutesD2DD00015F dataset. We applied a scaffold split to identify the ten most commonly occurring scaffolds within PaRoutes. From the clustering analysis, we selected 130 diverse target molecules to test our framework’s ability to replicate and improve upon patent-derived "gold-standard" routes.
  • Novelty assessment (ChemByDesign): To test generalization to unseen data, we selected 5 target molecules from ChemByDesign draghici2012chemistry. We specifically filtered for pathways published after 2021 with fewer than six steps. This filtering ensured that the ground-truth reactions were not present in our USPTO training data (which was last updated in September 2016), providing a test of zero-shot pathway prediction.
  • Hybrid comparative set: To benchmark against existing hybrid planners, we selected specific complex targets (e.g., dronabinol, arformoterol, 4-ethenyl-2-fluorophenol) used in prior studies by Levin et al. levin2022merging and RetroBioCat finnigan2021retrobiocat. These molecules were manually curated to allow for a direct qualitative comparison of pathway length, cost, and enzyme usage.
2.2   Model development
@@
We utilize the Modern Hopfield Network ramsauer2020hopfield based template prioritization architecture for our study. This architecture, initially introduced by Seidl et al. seidl2022improving, consists of three main components: a molecule encoder, a reaction template encoder, and one or more stacked or parallel Hopfield layers.
@@
The molecule encoder function, denoted as $h_{w}^{m}(\cdot)$, learns a relevant representation for the input molecule $m$. We utilize a fingerprint-based approach, specifically the extended connectivity fingerprint (ECFP), coupled with a fully connected neural network with weights $w$. This encoder maps a molecule to a dense representation $\mathbf{m}_{h}=h_{w}^{m}(m)$ of dimension $d_{m}$.
@@
Similarly, the reaction template encoder function, denoted as $h_{v}^{t}(\cdot)$, learns relevant representations of reaction templates. We employ a fully connected neural network with RDKit template fingerprints as input. The function is applied to the set of all templates $\mathcal{T}$, and the resulting vectors are concatenated column-wise into a matrix $\mathbf{T}_{h}=h_{v}^{t}(\mathcal{T})$ with shape $(d_{t},K)$, where $K$ is the number of templates. Both encoders utilize Xavier Initialization pmlr-v9-glorot10a and Batch Normalization ioffe2015batch to improve convergence and prevent vanishing gradients.
The core of the model consists of Hopfield layers, denoted as $g(\cdot,\cdot)$, which associate the molecule with the memory of templates. To perform the retrieval, the encoded molecule $\mathbf{m}_{h}$ and the template matrix $\mathbf{T}_{h}$ are projected into a common associative space to form the state pattern $\boldsymbol{\xi}$ and the stored patterns $\mathbf{X}$, respectively. The Hopfield layer then updates the molecule representation via an attention-like mechanism. The update rule yielding the new state $\boldsymbol{\xi}_{\text{new}}$ is defined as:
@@
$\boldsymbol{\xi}_{\text{new}}=\mathbf{X}\mathbf{p}=\mathbf{X}\cdot\mathrm{softmax}\left(\beta\mathbf{X}^{\top}\boldsymbol{\xi}\right)$ (1)
where $\beta$ is a learnable scaling parameter (inverse temperature), $\mathbf{p}$ is the vector of associations (probabilities) over the templates, and $\mathrm{softmax}$ is applied column-wise.
For the loss function, given a training pair $(m,t)$ and the set of all templates $\mathcal{T}$, the model aims to maximize the probability assigned to the correct template $t$. We employ the negative log-likelihood as the loss function, optimizing parameters using stochastic gradient descent via the AdamW optimizer. To prevent overfitting, we apply dropout and L2 regularization. Additionally, a post-processing fingerprint-based substructure screen is used during inference to filter out chemically non-applicable templates.
We train two sets of models on our datasets: the synthetic template prioritizer (consisting of five models) and the enzymatic template prioritizer. We utilized PyTorch pytorch2019 version 1.9.0 and Python version 3.8 to implement the model. These models take a target molecule as input and rank the most applicable rules within the collected dataset. Given the same inputs, the outputs from the five synthetic template prioritizer are taken individually, and a final ranking of the templates is collated using the highest predicted score. Furthermore, to fine-tune the models for optimal performance, hyperparameter tuning was performed using a one-factor-at-a-time (OFAT) approach. We tuned number of epochs, concatenation threshold, dropout rate, learning rate and some hopfield parameters, more information on the parameters and the values chosen can be found in Supplementary Method 1, 2 and Supplementary Table 1, 2. Each experiment involved modifying a single hyperparameter from a baseline configuration while keeping all other settings constant. The performance of each model was assessed using an evaluation score, with lower validation loss indicating better performance.
2.3   Searching algorithm